News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 6 October 2000 3.30pm EDT

During the last week, Chandra completed the observing schedule as planned.

Real-time procedures were completed during the last week to re-insert the LETG following the completion of the target of opportunity GB000926 at 4.30pm EDT on 29 Sept, retract the LETG as planned at 12.30pm EDT on 2 October, and to perform a test to determine the relative positions of the LETG bumper and retract limit switch. The bumper test was conducted at 12.30pm EDT on 4 October and the test data are being analyzed.

An unexpected Solid State Recorder (SSR) event occurred at 12:14pm on 29 Sept when recording switched from SSR-A to SSR-B without a commanded action. The switch occurred coincident with a commanded play-back from SSR-A and a fault tree analysis is under development. Following the switch, recording was continued on SSR-A, all data was recovered and all subsequent recording and playbacks have been nominal.

The planned observations for the coming week are as follows.

------------------------------------------
GRB000926              ACIS-S        Oct  8
Radiation Belt                       Oct  9
GRO J0422+32           ACIS-I        Oct 10
A520                   ACIS-I
B2 0738+313            ACIS-S
Abell 963              ACIS-S
3C 219                 ACIS-S        Oct 11
APM08279+5255          ACIS-S
NGC 507                ACIS-S
M31 (multiple)         HRC-I
Radiation Belt                       Oct 12
NGC 720                ACIS-S
A521                   ACIS-S        Oct 13
SMC X-1                ACIS-S
A133                   ACIS-S
HE0230-2130            ACIS-S        Oct 14
RX J0925.7-4758        ACIS-S
SMC X-1                ACIS-S
NGC 7582               ACIS-S
Radiation Belt
NGC 7582               ACIS-S
G11.2-0.3              ACIS-S        Oct 15
G29.7-0.3              ACIS-S

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports